OpenStudy (anonymous):

What are the solutions? 16x2 + 9 = 0 (1 point)

OpenStudy (anonymous):

@jdoe0001 help me please?

OpenStudy (anonymous):

@agent0smith

OpenStudy (jdoe0001):

\(\bf 16x^2 + 9 = 0\implies 16x^2 =-9\implies x^2= -\cfrac{9}{16}\implies x=\pm\sqrt{-\cfrac{9}{16}}\)

OpenStudy (agent0smith):

Subtract 9 from both sides first. .. Have you learned complex/imaginary numbers yet?

OpenStudy (jdoe0001):

\(\bf x=\pm\sqrt{-\cfrac{9}{16}}\implies x=\pm\sqrt{\cfrac{9}{16}}\cdot \sqrt{-1}\implies x =\sqrt{\cfrac{3^2}{4^2}}\cdot \sqrt{-1}\)

OpenStudy (anonymous):

these are my answer options, would it be the last one?